OpenStudy (chrisneedshelp):

how would 0.0044 / 24.1 = 0.00018??? could it be ratios? 2:1? i need help!

OpenStudy (anonymous):

That is simply the quotient of that ratio rounded to 2 places.

OpenStudy (anonymous):

use a calculator

OpenStudy (chrisneedshelp):

i did use a calculaor, 0.0044 / 24.1 = 1.82572 according to my laptops calculator?

OpenStudy (anonymous):

Do you want to express it as a rational number?

OpenStudy (chrisneedshelp):

im not quite sure, my teacher assisted me and told me that 0.0044 / 24.1 = 1.82 is wrong that it should be 0.0044 / 24.1 = 0.00018, how did he even get that? i think im supposed to apply ratios to this, but I do not know how to do a 2:1 ratio,

OpenStudy (anonymous):

As a ratio, 0.0044 / 24.1 = 44 : 241000, which can be reduced if you prefer.

OpenStudy (anonymous):

If you got something like "0.0044 / 24.1 = 1.82572" then it's possible that the answer was expressed in scientific notation. Did you see a "x10^-4" or "E-4" notation at the end of it?

OpenStudy (anonymous):

1.8 x10^-4 would be the correct expression in scientific notation using the correct number of significant digits. I still don't fully understand the question, though.

OpenStudy (chrisneedshelp):

yes! theres E-4 in the number!

OpenStudy (anonymous):

Then there ya' go. 1.8E-4 = 0.00018.

OpenStudy (chrisneedshelp):

how do i do that? (1.8E-4)?

OpenStudy (chrisneedshelp):

hold on ill upload some pictures

OpenStudy (chrisneedshelp):